Consider the given Probability Distribution. Then select all true statement/s.

XP(X)

--------------------------------------------------------------

5|0.27

6|0.23

7|0.23

8|0.17

9|0.10

10|0.00

Compute the expected value.

SELECT ALL APPLICABLE CHOICES

A)μ=2.7

B)μ=6.6

C)μ=3.4

D) None of These
Answer #1

Solution

Answer : B) 6.6

\mathbf{X} PX) X*P(X)
5 0.27 1.35
6 0.23 1.38
7 0.23 1.61
8 0.17 1.36
9 0.10 0.90
10 0.00 0.00
ΣΡ (X) = 1 X*P(X) = 6.6

Expected value = X*P(X) = 6.6